OpenStudy (hpaul24):

Find the sum of the infinite series 1 + 2 + 4 + 8 +…if it exists. 32,767 65,535 131,071 none

satellite73 (satellite73):

numbers get bigger and bigger right?

satellite73 (satellite73):

therefore so does the sum

geerky42 (geerky42):

inb4 \(-\dfrac1{12}\)

satellite73 (satellite73):

??

geerky42 (geerky42):

http://www.urbandictionary.com/define.php?term=inb4&defid=3632151 Since \(-\dfrac1{12}\) is Ramanujan sum, and this "math astounding fact" is so common, it's annoying. Just little tease, sorry

OpenStudy (kainui):

\[S=1+2+4+8+\cdots\]\[2\times S=2+4+8+16+\cdots\]\[1+2S=1+2+4+8+\cdots\]\[1+2S=S\]\[1+S=0\]\[S=-1\]

jimthompson5910 (jim_thompson5910):

@Kainui the proof falls apart when you go from `1+2S = S` to `1+S = 0`. Subtracting S from both sides has you using the form `infinity - infinity` but that's indeterminate.

geerky42 (geerky42):

@jim_thompson5910 I think he was just continuing the discussion of Ramanujan sum, which in that case, it's technically valid. It's just a way to assign value to divergent series.